Part A: What is the pH of a buffer prepared by adding 0.708 mol of the weak acid HA to 0.406 mol of NaA in 2.00 L of solution? The dissociation constant Ka of HA is 5.66×10−7.

Part B:What is the pH after 0.150 mol of HCl is added to the buffer from Part A? Assume no volume change on the addition of the acid.

Part C: What is the pH after 0.195 mol of NaOH is added to the buffer from Part A? Assume no volume change on the addition of the base.
Answer #1

- [H] CAT Parta: Acid buffer of weak acid HA and Naa Weak acid dissociated and attained as- HA => H+ + A, Ka=b! - [HA PH of tNOW DH = bko +100 [naA] = 6.25 + log 0-256 [PH = 5.73 = 5.73 After aditition at 0.150 mol of Hal to the butter pH becomes 5.7If you like the answer please give a thumbs up ?.